OpenStudy (anonymous):

hahah is this problem un solveable no one can solve it on here jeeze 2(x+4)^1/2+3(x+4)^3/2+(x+4)^5/2 factor using subsitution

OpenStudy (anonymous):

let u=x+4 2u^(1/2)+3u^(3/2)+u^(5/2) u^(1/2)(2+3u+u^2) u^(1/2)(u+1)(u+2) plug u=x+4 back.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Fan me for solving the unsolvable? :D

OpenStudy (anonymous):

what wud the final answer be ? x^(1/2)+2 (x+5)(x+6)?

OpenStudy (anonymous):

(x+4)^(1/2)(x+5)(x+7)

OpenStudy (anonymous):

(x+4)^(1/2)(x+5)(x+6). Oops.

OpenStudy (anonymous):

thanks